OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 36 y = x - 6

OpenStudy (zale101):

the second equation, pick a number for x and the results will be for why, and from their. Keep doing that until u have like 3 ordered pairs and draw a line for it . y = x - 6

OpenStudy (zale101):

for the first one, change the equation and solve for y. x2 + y2 = 36 change it to y^2 =36 + x^2

OpenStudy (anonymous):

how do i do it with the exponent

OpenStudy (shamil98):

x^2 + y^2 = 36 square everything x + y = 6

OpenStudy (shamil98):

x + y = 6 y = x - 6 x+ (x-6) = 6 2x - 6 = 6 2x = 12 x = 6 substitute that into y = x- 6 to get the value for y :)

OpenStudy (zale101):

i think i did something wrong... Does the lines intersect,so the y-value will be the same for each?

OpenStudy (anonymous):

thank you!

OpenStudy (shamil98):

zale , x^2 + y^2 = 36 is the same as x + y = 6

OpenStudy (shamil98):

and you did y^2 = 36 + x^2 it would be y^2 = 36 - x^2

OpenStudy (zale101):

oh, thx for the correction

OpenStudy (anonymous):

Shamil98...INCORRECT! You said that x^2 + y^2 = 36 is the same as x + y = 6. That is FALSE!!

OpenStudy (shamil98):

-_- if you square it then yes ..

OpenStudy (shamil98):

x + y = 6 is simplified form

OpenStudy (anonymous):

Arishagee......watch

OpenStudy (anonymous):

Shamil...NO!

OpenStudy (shamil98):

x = 6 , y = 0

OpenStudy (anonymous):

Shamil..last time..you are giving incorrect statements!

OpenStudy (anonymous):

Arishagee...you there?

OpenStudy (zale101):

@Easyaspi314 explain how's he wrong?

OpenStudy (shamil98):

omg, will you leave it? the second pair of solutions is (0,6) if thats what ur lookin for

OpenStudy (anonymous):

To solve that original syatem of equations: First note that we take y = x - 6 and substiutute that into y in the first equation, So the FIRST equation becomes x^2 + (x-6)^2 = 36 Simplify the above, you get x^2 + x^2 - 12x + 36 = 36 2x^2 - 12 x = 0 2x(x - 6) = 0 x = 0 or x = 6 If x = 0, y = -6; if x = 6, y = 0 Final solutions: x = 0 and y = -6 AND x = 6 and y= 0

OpenStudy (shamil98):

-6* yes i forgot the negative in the secoond pair

OpenStudy (shamil98):

what did i do wrong?.

OpenStudy (anonymous):

Shai...what you did was wrong..you CANNOT make a statement that x^2 + y^2 = 36 is the same as x + y = 6!!

OpenStudy (shamil98):

the question is closed, leave it . I made a false statement, okay? My bad? Jeez

OpenStudy (anonymous):

Thats nonsense. The x^2 + y^2 = 36 is a circle whose radius is 6 and center is at the origin, while x+ y = 6 is a straight line. Two different animals!

OpenStudy (anonymous):

Shamil..I must and needed to correct you, as others may walk away assuming that if they see x^2 + y^2 = 36 it would be the same as x+ y = 6, and it is NOT the same. Besides, I have two pairs of solutions that work correctly. I did it mathematically correct.

OpenStudy (anonymous):

The algebraic way: Use substitution: \[\begin{split} x^2+y^2&=36\\ x^2+(x-6)^2&=36&\quad &y=x-6\\ \end{split} \]Now simplify you have a quadratic function in the ends.

OpenStudy (anonymous):

wio..thats what I just showed them

OpenStudy (zale101):

that was the easiest method ever!

OpenStudy (zale101):

the second equation was y=x-6, i should've plugged it to y of the first equation -.-

OpenStudy (anonymous):

Anyway:\[ (x+y)^2=6^2\\ x^2+2xy+y^2=36 \]Using elimination you get:\[ 2xy=0 \]Meaning there are solutions when \(x=0\) and \(y=0\).

OpenStudy (anonymous):

Zale...Good Morning...thanks for smelling the coffee.

OpenStudy (shamil98):

I showed a simple method, I made an incorrect statement, leave it at that.

OpenStudy (zale101):

keep convos in pm >.>

OpenStudy (anonymous):

Why you get 2 medals?

OpenStudy (shamil98):

Zale and the asker gave me one each lol

OpenStudy (zale101):

shamil, ur answer was explained perfectly, that's why i gave u that medal xD

OpenStudy (anonymous):

I figured out a way to do it with elimination and substitution. That is pretty boss.

OpenStudy (shamil98):

I did it by making it simple :D

OpenStudy (anonymous):

You didn't do it right though. You were wrong.

OpenStudy (shamil98):

The solutions were right ._.

OpenStudy (zale101):

-.-

OpenStudy (zale101):

i'm keep getting notifications :\

OpenStudy (zale101):

i*

OpenStudy (shamil98):

yep >.>

OpenStudy (zale101):

Close this question @arishagee, it's solved and done

OpenStudy (shamil98):

it is closed o.o